# Thermodynamics of Modified Chaplygin Jacobi And Modified Chaplygin Abel Gas: Stability Analysis and Observational Constraints

### v1

@@ 3   Thermodynamic Analysis of Modified Chaplygin Abel Gas
$\rho=\left[\dfrac{Bc^{2}}{(A+1)\left(c^{2}+e^{2}\left(1+\frac{1}{V^{e^{2}(A+1)(\alpha+1)}c^{2}\left(-\frac{B}{2D}\right)^{e^{2}(A+1)}-1}\right)\right)}\right]^{\frac{1}{\alpha+1}}$ (59)
@@
$\rho=\left[\frac{B}{(A+1)}\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)\right]^{\dfrac{1}{\alpha+1}}$ (60)
@@
$V_{0}^{(1+\alpha)}=\dfrac{2D}{B}$
. From the equation (60) it is clearly obtainable that at very low volume i.e. for $V\to 0$, the energy density becomes very high. Alternatively, at very large volume i.e for $V\to\infty$ the energy represents the mixture of two fluids with one has constant energy density and other has a volume dependent energy density given as,
$\rho\approx\left(\frac{B}{A+1}\right)^{\frac{1}{1+\alpha}}+\frac{1}{1+\alpha}\left(\frac{e^{2}}{c^{2}}\right)\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}$ (61)
@@
$p=-\dfrac{Bc^{2}}{\left[\frac{B}{A+1}\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)\right]^{\frac{\alpha}{\alpha+1}}}$ (62)
@@
$p\approx-Bc^{2}\left[\left(\frac{B}{A+1}\right)^{-\frac{\alpha}{1+\alpha}}\left(\frac{V}{V_{0}}\right)^{e^{2}\alpha(1+A)}\right]\approx 0$ (63)
@@
$p\approx-Bc^{2}\left[\frac{B}{A+1}\right]^{-\frac{\alpha}{\alpha+1}}$ (64)
@@ 3.1   EoS Parameter
The equation of state parameter is expressed as $\omega=\dfrac{p}{\rho}$. By substituting the values of pressure and energy density, the resulting equation of state parameter for the Modified Chaplygin Abel Gas is given by:
$\omega=-\frac{Bc^{2}}{\left[\frac{B\left(c^{2}+e^{2}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{c^{2}(A+1)}\right]^{\frac{1}{\alpha+1}}\left[\left(\frac{B\left(c^{2}+e^{2}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{c^{2}(A+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\alpha}}$ (65)
It represents a very small constant value of EoS parameter at small volume ($V\to 0$) and at large volume ($V\to\infty$) the EoS parameter becomes negative and volume dependent and is plotted in fig. 6(a).
@@ 3.2   Decelerating Parameter
The deceleration parameter $q$ indicates the rate of expansion of the universe. A negative value of $q$ signifies that the universe is undergoing an accelerated expansion phase, while a positive value indicates a decelerating phase of expansion.
Using the definition from equation (17), we obtained
$q=-\frac{3Bc^{2}}{2\left[\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right]^{\frac{1}{\alpha+1}}\left[\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\alpha}}+0.5$ (66)
@@
Figure 6: Graphical representation illustrating the relationship between the volume $V$ and two key cosmological indicators in the Modified Chaplygin Abel gas (MCAG) framework: the equation of state coefficient $\omega$ and the deceleration factor $q$.
3.3   Square of The Speed of Sound
@@
By using the definition from MCJG we have obtained,
$v_{s}^{2}=\frac{B\alpha c^{2}}{\left[\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right]^{\frac{1}{\alpha+1}}\left[\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\alpha}}+(A+1)\left(2c^{2}+e^{2}\right)-1$ (67)
@@ 3.4   Thermodynamic Stability
$\left(\frac{\partial p}{\partial V}\right)_{T}=-\frac{B\alpha e^{4}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}(A+1)}{V\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right)^{\frac{\alpha}{\alpha+1}}}$ (68)
From the above expression it is readily inferred that $B\neq 0$, because if so then $\left(\dfrac{\partial p}{\partial V}\right)_{T}=0$ everywhere irrespective of the volume and other parameters. The figure. 7(b) shows that the pressure reduces as the universe expands adiabatically. This imposes that the pressure of MCJG model defined in (55) can attain any value without any constrain.
@@
$T=\dfrac{\partial U}{\partial S}$ (69)
$=\Big{(}\dfrac{\partial U}{\partial D}\Big{)}\Big{(}\dfrac{\partial D}{\partial S}\Big{)}$
When examining the thermodynamic properties of the system, a critical consideration emerges regarding the parameters $A$ and $D$. Should these parameters be considered entropy-independent constants, the system exhibits a problematic behavior wherein the temperature approaches zero irrespective of the gas’s pressure or volumetric state. This scenario creates a thermodynamically inconsistent situation where the zero-temperature isotherm ($T=0$) becomes coincident with an isentropic pathway ($S=\text{const}$). Such behavior fundamentally conflicts with established thermodynamic principles, particularly the third law. Consequently, a rigorous analysis of the Modified Chaplygin Abel Gas (MCAG) thermodynamic stability requires the postulation that entropy dependence exists in at least one of these parameters $S$.
Now, the internal energy $U$ is given as
$U=V\left(\frac{Bc^{2}}{(A+1)\left(c^{2}+e^{2}\left(1+\frac{1}{V^{e^{2}(A+1)(\alpha+1)}c^{2}\left(-\frac{B}{2D}\right)^{e^{2}(A+1)}-1}\right)\right)}\right)^{\frac{1}{\alpha+1}}$ (70)
@@
$\frac{\partial D}{\partial S}=\frac{\nu^{2\alpha+2}(S\tau)^{-\alpha-1}(-\alpha-1)}{S}$ (72)
@@
$-F^{2}OYN\left(E\left(-Z+XEY+XOY\right)+\left(Z-XEY\right)(E+O)\right)$
@@
$X$ $=V^{e^{2}(A+1)(\alpha+1)},$
$Y$ $=(-B\nu^{-2\alpha-2}(S\tau)^{\alpha+1})^{e^{2}(A+1)},$
$Z$ $=2^{e^{2}(A+1)},$
$E$ $=c^{2},$
$F$ $=V^{e^{2}(A+1)(\alpha+1)+1},$
$N$ $=(A+1)(\alpha+1),$
$O$ $=e^{2}.$
@@
$T=\left(\dfrac{1}{S}\right)\left[\frac{-Bc^{2}}{\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{A+1}\right)^{\frac{\alpha}{\alpha+1}}}+\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{A+1}\right)^{\frac{1}{\alpha+1}}\right]$ (74)
@@ 3.5   Reduced Parameters
$p(T)=-\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}$ (75)
Here, $T^{*}$ represents an integration constant, and the condition $0<T<T^{*}$ indicates that $T^{*}$ is the highest temperature achievable by the gas. Now, let us consider the initial values of this system to be $V=V_{i}$, $p=p_{i}$, $\rho=\rho_{i}$, and $T=T_{i}$.
Now, from the equation (60), we can find the expression of the arbitrary parameter $D$ in terms of the initial values as,
$\left(\frac{2D}{B}\right)^{e^{2}(1+A)}=\left[\left(\frac{A+1}{B}\right)\rho_{i}^{1+\alpha}-1\right]\left(\frac{c}{e}\right)^{2}V_{i}^{e^{2}(1+A)(1+\alpha)}$ (76)
@@
$\rho=\rho_{i}\left[\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{e^{2}(1+A)(1+\alpha)}\right]^{\frac{1}{1+\alpha}}$ (77)
$p=-Bc^{2}\left[\rho_{i}\left(\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{e^{2}(1+A)(1+\alpha)}\right)^{\frac{1}{1+\alpha}}\right]^{-\alpha}$ (78)
@@
$\epsilon$ $=$ $\dfrac{\rho}{\rho_{i}}$ (79)
$v$ $=$ $\dfrac{V}{V_{i}}$ (80)
$P$ $=$ $\dfrac{p}{\left(\frac{B}{A+1}\right)^{-\frac{\alpha}{1+\alpha}}}$ (81)
$\gamma$ $=$ $\dfrac{B}{(A+1)\rho_{i}^{1+\alpha}}$ (82)
$t$ $=$ $\dfrac{T}{T_{i}}$ (83)
$t^{*}$ $=$ $\dfrac{T^{*}}{T_{i}}$ (84)
@@
$\epsilon$ $=\left[\gamma+\dfrac{1-\gamma}{v^{e^{2}(1+A)(1+\alpha)}}\right]^{\frac{1}{1+\alpha}}$ (85)
$P$ $=-Bc^{2}\gamma^{\frac{\alpha}{1+\alpha}}\cdot\left[\gamma+\dfrac{1-\gamma}{v^{e^{2}(1+A)(1+\alpha)}}\right]^{-\frac{\alpha}{1+\alpha}}$ (86)
$p(T)$ $=-\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{t}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}$ (87)
@@
$\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{1}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}=Bc^{2}\gamma^{\dfrac{\alpha}{\alpha+1}}$ (88)
@@
$t^{*}=\left(\dfrac{1}{Bc^{2}}\right)^{\frac{1}{1+\alpha}}\left[\dfrac{1}{1-\gamma}\right]^{\frac{\alpha}{\alpha+1}}$ (89)
@@
$\gamma=1-(Bc^{2})^{-\frac{1}{\alpha}}\left(t^{*}\right)^{-\frac{\alpha+1}{\alpha}}$ (90)
Our analysis establishes two crucial temperature benchmarks for the Modified Chaplygin Abel Gas: an upper bound temperature $T^{*}$, which reaches $10^{32}$ units corresponding to conditions during the Planck epoch, and the current cosmic background temperature $T_{i}$, measured at $2.7$ units. These boundary conditions enable us to perform the following derivation:
$\gamma=1-(Bc^{2})^{-\frac{1}{\alpha}}\left(10^{32}\right)^{-\frac{\alpha+1}{\alpha}}\approx 1$ (91)
In this scenario, the energy density of the universe composed of Modified Chaplygin Abel Gas (MCAG) tends toward its limiting value expressed as $\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}$.
The internal energy of the system becomes,
$U=\dfrac{(Bc^{2})^{\frac{1}{\alpha}}V}{\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}$ (92)
Now, using the thermodynamic relation (53),
we obtain,
$\dfrac{(Bc^{2})^{\frac{1}{\alpha}}}{\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}=\dfrac{Bc^{2}}{\rho^{\alpha}}\left[\frac{T}{p}\frac{dp}{dT}-1\right]$ (93)

### v2

@@ 3   Thermodynamic Analysis of Modified Chaplygin-Abel Gas
$\rho=\left[\dfrac{Bc^{2}}{(A+1)\left(c^{2}+e^{2}\left(1+\frac{1}{V^{e^{2}(A+1)(\alpha+1)}c^{2}\left(-\frac{B}{2D}\right)^{e^{2}(A+1)}-1}\right)\right)}\right]^{\frac{1}{\alpha+1}}.$ (51)
@@
$\rho=\left[\frac{B}{(A+1)}\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)\right]^{\dfrac{1}{\alpha+1}}.$ (52)
@@
$V_{0}^{(1+\alpha)}=\dfrac{2D}{B}\quad.$
From the equation (52) it is clearly obtainable that at very low volume i.e. for $V\to 0$, the energy density becomes very high. Alternatively, at very large volume i.e for $V\to\infty$ the energy represents the mixture of two fluids with one has constant energy density and other has a volume dependent energy density given as,
$\rho\approx\left(\frac{B}{A+1}\right)^{\frac{1}{1+\alpha}}+\frac{1}{1+\alpha}\left(\frac{e^{2}}{c^{2}}\right)\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}.$ (53)
@@
$p=-\dfrac{Bc^{2}}{\left[\frac{B}{A+1}\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)\right]^{\frac{\alpha}{\alpha+1}}}\quad.$ (54)
@@
$p\approx-Bc^{2}\left[\left(\frac{B}{A+1}\right)^{-\frac{\alpha}{1+\alpha}}\left(\frac{V}{V_{0}}\right)^{e^{2}\alpha(1+A)}\right]\approx 0,$ (55)
@@
$p\approx-Bc^{2}\left[\frac{B}{A+1}\right]^{-\frac{\alpha}{\alpha+1}}.$ (56)
@@ 3.1   EoS Parameter
The equation of state parameter is expressed as $\omega=\dfrac{p}{\rho}$. By substituting the values of pressure and energy density, the resulting equation of state parameter for the Modified Chaplygin-Abel Gas is given by:
$\omega=-\frac{Bc^{2}}{\left[\frac{B\left(c^{2}+e^{2}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{c^{2}(A+1)}\right]^{\frac{1}{\alpha+1}}\left[\left(\frac{B\left(c^{2}+e^{2}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{c^{2}(A+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\alpha}}\quad.$ (57)
It represents a very small constant value of EoS parameter at small volume ($V\to 0$) and at large volume ($V\to\infty$) the EoS parameter becomes negative and volume dependent and is plotted in fig. 6(a). It is worthy to note that for the chosen values of $B=c=1$ and $V_{0}=A=0$, the EoS $\omega$ becomes $-1$ which is the dark energy equation of state parameter in the standard $\Lambda$CDM paradigm irrespective of the values of $\alpha$ and $e$.
@@ 3.2   Decelerating Parameter
The deceleration parameter $q$ indicates the rate of expansion of the universe. A negative value of $q$ signifies that the universe is undergoing an accelerated expansion phase, while a positive value indicates a decelerating phase of expansion. Using the definition from equation (14), we obtained
$q=-\frac{3Bc^{2}}{2\left[\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right]^{\frac{1}{\alpha+1}}\left[\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\alpha}}+0.5\quad.$ (58)
@@
Figure 6: Graphical representation illustrating the relationship between the volume $V$ and two key cosmological indicators in the Modified Chaplygin-Abel gas framework: the equation of state coefficient $\omega$ and the deceleration factor $q$.
3.3   Square Speed of Sound
@@
By using the definition from subsect. 2.3 we have obtained,
$v_{s}^{2}=\frac{B\alpha c^{2}}{\left[\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right]^{\frac{1}{\alpha+1}}\left[\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\alpha}}+(A+1)\left(2c^{2}+e^{2}\right)-1\quad.$ (59)
@@ 3.4   Thermodynamic Stability
$\left(\frac{\partial p}{\partial V}\right)_{T}=-\frac{B\alpha e^{4}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}(A+1)}{V\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{(A+1)}\right)^{\frac{\alpha}{\alpha+1}}}\quad.$ (60)
From the above expression it is readily inferred that $B\neq 0$, because if so then $\left(\dfrac{\partial p}{\partial V}\right)_{T}=0$ everywhere irrespective of the volume and other parameters. The figure. 7(b) shows that the pressure reduces as the universe expands adiabatically. This imposes that the pressure of MCAG model defined in (47) can attain any value without any constrain.
@@
$T=\dfrac{\partial U}{\partial S}=\Big{(}\dfrac{\partial U}{\partial D}\Big{)}\Big{(}\dfrac{\partial D}{\partial S}\Big{)}.$ (61)
When examining the thermodynamic properties of the system, a critical consideration emerges regarding the parameters $A$ and $D$. Should these parameters be considered entropy-independent constants, the system exhibits a problematic behavior wherein the temperature approaches zero irrespective of the gas’s pressure or volumetric state. This scenario creates a thermodynamically inconsistent situation where the zero-temperature isotherm ($T=0$) becomes coincident with an isentropic pathway ($S=\text{const}$). Such behavior fundamentally conflicts with established thermodynamic principles, particularly the third law. Consequently, a rigorous analysis of the MCAG thermodynamic stability requires the postulation that entropy dependence exists in at least one of these parameters $S$. Now, the internal energy $U$ is given as
$U=V\left(\frac{Bc^{2}}{(A+1)\left(c^{2}+e^{2}\left(1+\frac{1}{V^{e^{2}(A+1)(\alpha+1)}c^{2}\left(-\frac{B}{2D}\right)^{e^{2}(A+1)}-1}\right)\right)}\right)^{\frac{1}{\alpha+1}}.$ (62)
@@
$\frac{\partial D}{\partial S}=\frac{\nu^{2\alpha+2}(S\tau)^{-\alpha-1}(-\alpha-1)}{S}\quad.$ (64)
@@
$-F^{2}OYN\left(E\left(-Z+XEY+XOY\right)+\left(Z-XEY\right)(E+O)\right)\quad,$
@@
$X$ $=V^{e^{2}(A+1)(\alpha+1)},\quad Y=(-B\nu^{-2\alpha-2}(S\tau)^{\alpha+1})^{e^{2}(A+1)},\quad Z=2^{e^{2}(A+1)},$
$E$ $=c^{2},\quad F=V^{e^{2}(A+1)(\alpha+1)+1},\quad N=(A+1)(\alpha+1),\quad O=e^{2}.$
@@
$T=\left(\dfrac{1}{S}\right)\left[\frac{-Bc^{2}}{\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{A+1}\right)^{\frac{\alpha}{\alpha+1}}}+\left(\frac{B\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)}{A+1}\right)^{\frac{1}{\alpha+1}}\right]\quad.$ (66)
@@ 3.5   Reduced Parameters
$p(T)=-\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}.$ (67)
Here, $T^{*}$ represents an integration constant, and the condition $0<T<T^{*}$ indicates that $T^{*}$ is the highest temperature achievable by the gas. Now, let us consider the initial values of this system to be $V=V_{i}$, $p=p_{i}$, $\rho=\rho_{i}$, and $T=T_{i}$. Now, from the equation (52), we can find the expression of the arbitrary parameter $D$ in terms of the initial values as,
$\left(\frac{2D}{B}\right)^{e^{2}(1+A)}=\left[\left(\frac{A+1}{B}\right)\rho_{i}^{1+\alpha}-1\right]\left(\frac{c}{e}\right)^{2}V_{i}^{e^{2}(1+A)(1+\alpha)}.$ (68)
@@
$\rho=\rho_{i}\left[\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{e^{2}(1+A)(1+\alpha)}\right]^{\frac{1}{1+\alpha}},$ (69)
$p=-Bc^{2}\left[\rho_{i}\left(\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{e^{2}(1+A)(1+\alpha)}\right)^{\frac{1}{1+\alpha}}\right]^{-\alpha}.$ (70)
@@
$\epsilon=\dfrac{\rho}{\rho_{i}},\quad v=\dfrac{V}{V_{i}},\quad P=\dfrac{p}{\left(\frac{B}{A+1}\right)^{-\frac{\alpha}{1+\alpha}}},\quad\gamma=\dfrac{B}{(A+1)\rho_{i}^{1+\alpha}},\quad t=\dfrac{T}{T_{i}},\quad t^{*}=\dfrac{T^{*}}{T_{i}}.$ (71)
@@
$\epsilon$ $=\left[\gamma+\dfrac{1-\gamma}{v^{e^{2}(1+A)(1+\alpha)}}\right]^{\frac{1}{1+\alpha}},$ (72)
$P$ $=-Bc^{2}\gamma^{\frac{\alpha}{1+\alpha}}\cdot\left[\gamma+\dfrac{1-\gamma}{v^{e^{2}(1+A)(1+\alpha)}}\right]^{-\frac{\alpha}{1+\alpha}},$ (73)
$p(T)$ $=-\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{t}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}.$ (74)
@@
$\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{1}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}=Bc^{2}\gamma^{\frac{\alpha}{\alpha+1}}.$ (75)
@@
$t^{*}=\left(\dfrac{1}{Bc^{2}}\right)^{\frac{1}{1+\alpha}}\left[\dfrac{1}{1-\gamma}\right]^{\frac{\alpha}{\alpha+1}},$ (76)
@@
$\gamma=1-(Bc^{2})^{-\frac{1}{\alpha}}\left(t^{*}\right)^{-\frac{\alpha+1}{\alpha}}.$ (77)
Our analysis establishes two crucial temperature benchmarks for the Modified Chaplygin-Abel Gas: an upper bound temperature $T^{*}$, which reaches $10^{32}$ units corresponding to conditions during the Planck epoch, and the current cosmic background temperature $T_{i}$, measured at $2.7$ units. These boundary conditions enable us to perform the following derivation:
$\gamma=1-(Bc^{2})^{-\frac{1}{\alpha}}\left(10^{32}\right)^{-\frac{\alpha+1}{\alpha}}\approx 1\quad.$ (78)
In this scenario, the energy density of the universe composed of MCAG tends toward its limiting value expressed as $\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}$. The internal energy of the system becomes,
$U=\dfrac{(Bc^{2})^{\frac{1}{\alpha}}V}{\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}\quad.$ (79)
Now, using the thermodynamic relation (45), we obtain,
$\dfrac{(Bc^{2})^{\frac{1}{\alpha}}}{\left[(Bc^{2})^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}=\dfrac{Bc^{2}}{\rho^{\alpha}}\left[\frac{T}{p}\frac{dp}{dT}-1\right]\quad.$ (80)